# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

### v1

@@ 6.1 Training Stage
Figure 2: For the training stage with active data collection, we report (a) the training loss vs. the training iterations, and (b) the evaluation accuracy vs. the training iterations.
@@
Table 2: Comparison of different methods in terms of evaluation accuracy and training time in seconds. Results are averaged over five runs.
6.2 Deployment Stage
In this part, we evaluate our method on deployment stage.
Contenders.
In the deployment stage, we compare with:
  • Random: A baseline strategy that randomly select responses from responses generated by the current policy model.
  • Best-Two: A baseline strategy that selects the best two responses from responses generated by the current policy model.
  • Iterative-RLHF (Dong et al., 2024): A method that takes the best and the worst responses from the current policy model to construct a new preference pair for RLHF iteratively.
Deployment Stage Analysis.
For the scenario of deployment stage, we evaluate on a pre-processed online-variant of the Ultrafeedback-binarized dataset, where we divide the dataset into 20 chunks and process them sequentially to simulate the online deployment scenario. As shown in Figure 3, our method consistently outperforms the baselines in terms of both average cumulative rewards and win rates. The average cumulative rewards in Figure 3(a) show that our method learns more effectively from interactions, achieving higher rewards over time. The win rates in Figure 3(b) demonstrate that our method makes better decisions when selecting between response pairs, outperforming both the baseline methods of Random and Best-Two, and also better than Iterative-RLHF which selects the best and worst response pairs for training. The superior performance of our method further validates the effectiveness of our theoretically grounded action selection strategy. This success stems from an efficient online update mechanism that enables rapid adaptation to new preference feedback, as well as the design of the promising set, which balances high-reward responses for immediate performance with sufficient exploration for future model improvements. These results align with our theoretical analysis in Section 5, validating the practical effectiveness of our proposed algorithm in the deployment stage.
(a) average cumulative rewards
(b) win rates
Figure 3: For the deployment stage, we divide the Ultrafeedback dataset into 20 chunks, and report (a) the average cumulative rewards and the (b) win rates of different methods.
7 Conclusion and Future Work
In this work, we present a unified framework of the RLHF pipeline from the view of contextual bandits, leveraging the Bradley-Terry preference model with a linearly parameterized reward function. We decompose the RLHF process into two stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase. We then develop novel algorithms for each stage, applying online mirror descent with a tailored local norm to replace standard maximum likelihood estimation, and designing distinct query criteria for different stages. Our theoretical analysis shows that our method improves upon existing methods in both statistical and computational efficiency. Finally, we train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our approach. The empirical results validate our theoretical findings and confirm the effectiveness of our proposed method.
@@ 7 Conclusion and Future Work
While our work advances the theoretical understanding of RLHF, several important directions remain for future work. First, we considered the fixed feature mapping $\phi$ for the reward model, but in practice, this mapping may evolve during the training process. Thus, analyzing this dynamic feature mapping is an important future work. Second, though we focused on the Bradley-Terry model, there are other preference models for human feedback, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975). Extending our results to these preference models is another promising direction.
@@ Appendix A Useful Lemmas
For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as
~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$
Then, for the following update rule
@@
where the last inequality is by $\|\widetilde{\theta}_{1}-\theta^{*}\|_{\mathcal{H}_{1}}^{2}\leq\lambda\|\widetilde{\theta}_{1}-\theta^{*}\|_{2}^{2}\leq 4\lambda B^{2}$. Set $\zeta=2\eta$ finishes the proof. ∎
@@ Appendix B Proof of Lemma 1
To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as
$[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}\quad\text{and}\quad[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)},$
where $[\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
@@
Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with
$[\sigma^{-1}(p)]=\log({q}/({1-q}))\quad\text{and}\quad[\sigma^{-1}(p)]_{0}=\log((1-q)/{q}).$
Then, we decompose the regret into two terms by introducing an intermediate term.
$\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})=$ $\underbrace{\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})}_{\mathtt{term~{}(b)}}$
@@
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$ $\leq 2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)$
$+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda L_{\theta}^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}.$
Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
$12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}\leq\left(12\sqrt{2}BL^{3}\eta-\frac{\lambda}{7}\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}\leq 0$
@@ Appendix C Proof of Theorem 1
Define $J(\pi)=\mathbb{E}_{x\sim\rho}[r\left(x,\pi(x)\right)]$, $J^{\prime}(\pi)=J(\pi)-\langle\theta^{*},v\rangle$, we have
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right).$
@@
$\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi_{T}(s))-v)\right]-\mathbb{E}_{x\sim\rho}\left[\theta^{*\top}(\phi(s,\pi_{T}(s))-v)\right]\leq 0,$ (19)
@@
$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)$
=\ = $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}(\phi(s,\pi^{*}(s))-v)\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi^{*}(s))-v)\right]$
=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
@@
Since it holds $\widetilde{\theta}_{T}\in\mathcal{C}_{T}$ with probability at least $1-\delta$ by Lemma 1, we have $\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$ and $\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$. Thus, we obtain
$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (20)
@@
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)$ $=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right)$
$\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$
$\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
@@ Appendix D Proof of Theorem 2
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}.$

### v2

@@ 7.2.1 Passive Data Collection
(c) evaluation accuracy and training time
Figure 5: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
(a) training loss
(b) evaluation accuracy
Method ACC (%) Time (s)
Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
(c) evaluation accuracy and training time
Figure 6: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
7.2.2 Active Data Collection
In this experimental setup, we constrain the algorithm to select only 6,400 samples from the full training dataset, based on different data selection strategies. We evaluate the performance of both the standard MLE-based method and our proposed OMD-based method under this limited data regime. To assess the effectiveness of the data selection strategy itself, we compare our approach against a random selection baseline. We report the results across four key metrics: (a) training loss curve, (b) evaluation accuracy curve and (c) final evaluation accuracy and training time. Figure 5 and Figure 6 demonstrate that our OMD-based method achieves competitive performance with the MLE-based method for both data collection strategies, while improving the training time by significantly. Moreover, our data selection strategy outperforms the random selection strategy, demonstrating that our method can effectively select informative data to improve the performance.
7.2.3 Deployment-Time Adaptation
To emulate real-world deployment scenarios, we partition the dataset into 20 sequential chunks and process them incrementally. This setup allows us to evaluate the adaptability and robustness of different action selection strategies over time. We benchmark our proposed selection strategy against three baselines: (i): random selection, (ii): the best and second best actions, and (iii): the best and worst actions. We combine the above strategies with MLE-based and OMD-based reward estimators. We report three key metrics: (a) the average cumulative reward for MLE-based methods, (b) the average cumulative reward for OMD-based methods, and (c) the win rate of OMD-based methods over MLE-based counterparts under identical selection strategies. As illustrated in Figure 7 and Figure 8, our action selection strategy consistently outperforms the baselines across both estimator types. These results highlight the strength of our approach in balancing the exploitation of high-reward responses with sufficient exploration to drive continual model improvement. Moreover, the win rate comparisons suggest that our OMD-based estimator shows competitive performance relative to its MLE-based counterpart, further validating its practical effectiveness.
(a) Reward of MLE-based methods
(b) Reward of OMD-based methods
(c) Win rates, all methods
Figure 7: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
(a) Reward of MLE-based methods
(b) Reward of OMD-based methods
(c) Win rates, all methods
Figure 8: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
8 Conclusion
In this work, we study the problem of online RLHF from the perspective of contextual bandits. To address the linear growth in computational complexity with respect to the number of iterations in reward modeling, we propose a novel one-pass reward modeling algorithm that achieves constant-time complexity per iteration. Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information. We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation, and design tailored algorithms for each scenario. We provide theoretical guarantees that our approach achieves superior statistical and computational efficiency relative to existing methods. Finally, we provide practical implementations of our method and conduct experiments using the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, empirically validating its practical effectiveness.
@@ 8 Conclusion
While our work advances both the statistical and computational understanding of online RLHF, several important directions remain for future exploration. First, we assume a fixed feature mapping for the reward model; however, in practice, this mapping may evolve throughout the training process. Analyzing the impact of such dynamically changing feature representations presents a compelling direction for future research. Second, although our analysis is based on the Bradley-Terry model, extending the framework to other preference models, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975), is another promising avenue that may broaden the applicability of our results.
@@ Appendix A Useful Lemmas
For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$ Then, for the following update rule
@@
where the last inequality is by $\|\widetilde{\theta}_{1}-\theta^{*}\|_{\mathcal{H}_{1}}^{2}\leq\lambda\|\widetilde{\theta}_{1}-\theta^{*}\|_{2}^{2}\leq 4\lambda B^{2}$. Set $\zeta=2\eta$ ends the proof. ∎
@@ Appendix B Proof of Lemma 1
To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as $[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}$ and $[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)}$, where $[\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
@@
Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with $[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))$ and $[\sigma^{-1}(p)]_{0}=\log((1-p)/{p})$. Then, we decompose the regret into two terms by introducing an intermediate term.
$\sum_{i=1}^{t}\ell_{i}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})=$ $\underbrace{\sum_{i=1}^{t}\ell_{i}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{i}(q_{i},y_{i})}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{t}\ell_{i}(q_{i},y_{i})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})}_{\mathtt{term~{}(b)}}$
@@
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$ $\leq 12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}+C.$
where $C=2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$. Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
$12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}$
[…]
@@ Appendix C Proof of Theorem 1
Define $J(\pi)=\mathbb{E}_{x\sim\rho}[r\left(x,\pi(x)\right)]$, we have
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\left(J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J\left(\pi_{T}\right)\right).$
@@
$\widetilde{J}\left(\pi_{T}\right)-J\left(\pi_{T}\right)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi_{T}(s))\right]-\mathbb{E}_{x\sim\rho}\left[\theta^{*\top}\phi(s,\pi_{T}(s))\right]\leq 0,$ (12)
@@
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)=$ $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}\phi(s,\pi^{*}(s))\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi^{*}(s))\right]$
=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\phi(x,\pi^{*}(x))\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
@@
Since it holds $\theta^{*}\in\mathcal{C}_{T}$ with probability at least $1-\delta$ by Lemma 1, we have $\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$ and $\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$. Thus, we obtain
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (13)
@@
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
@@ Appendix D Proof of Theorem 2
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)=\frac{1}{\kappa}V_{t}.$